OpenStudy (anonymous):

According to the table, if the interest rate on this account is 10 percent, how much money will you have in the account at the end of the fourth year? A. $161.05 B. $146.41 C. $133.10 D. $120.00

OpenStudy (anonymous):

http://awesomescreenshot.com/0c52i4xea2

OpenStudy (anonymous):

I got D. because of the pattern from the Principal End of the Year : $110.00 $121.00 $133.10 $146. 41 by seeing 10 +11 = 21 + 12 = 33 + 13 = 46

OpenStudy (anonymous):

So B. $146.41